OpenStudy (anonymous):

Pamela's age is two times Jiri's age. The sum of their ages is 81. What is Jiri's age?

OpenStudy (anonymous):

p = 3j p + j = 64 4j = 64 j = 16

OpenStudy (anonymous):

let us say that Pamela's age= P Jiri's age=J then P=2J and P+J=81 u can put 2J instead of P in the second equation 2J+J=81 3J=81 J=27

OpenStudy (anonymous):

j= jiri p=2j j+2j=81 3j=81 j=27 jiri is 27 and pamela is 54